Solve the problem.
two pipes can fill a large tank in 10 hours. One of the pipes, used alone, takes 15 hours longer than the other to fill the tank. How long would each pipe take to fill the tank alone?
A) 10 hours for one
20 hours for the other
B) 15 hours for one
25 hours for the other
c) 15 hours for one
30 hours for the other
d) 12.5 hours for one
25 hours for the other
E) 25 hours for one
50 hours for the other

Answer by lwsshak3(11628)

two pipes can fill a large tank in 10 hours. One of the pipes, used alone, takes 15 hours longer than the other to fill the tank. How long would each pipe take to fill the tank alone?
***
let x=hours the first pipe takes to fill the tank
Its rate=1/x
x+15=hours the second takes to fill the tank
Its rate=1/(x+15)
1/10=rate of both pipes working together
sum of individual rates = rate working together
..
1/x+1/(x+15)=1/10
LCD: x(x+15)
x+15+x=x(x+15)/10
2x+15=(x^2+15x)/10
20x+150=x^2+15x
x^2-5x-150=0
(x-15)(x+10)=0
x=-10(reject)
x=15
x+15=30
hours the first pipe takes to fill the tank=15
hours the second pipe takes to fill the tank=30
answer is c)